Solve the given quadratic equation 2x+ x + 1 = 0.

Grade:12

1 Answers

Harshit Singh
3 years ago
Welcome to AskIITians

Given quadratic equation: 2x^2+ x + 1 = 0

Now, compare the given quadratic equation with the general form ax^2+ bx + c = 0

On comparing, we get

a = 2, b = 1 and c = 1

Therefore, the discriminant of the equation is:

D = b^2– 4ac

Now, substitute the values in the above formula

D = (1)^2– 4(2)(1)

D = 1- 8

D = -7

Therefore, the required solution for the given quadratic equation is

x =[-b ± √D]/2a

x = [-1 ± √-7]/2(2)

We know that, √-1 = i

x = [-1 ± √7i] / 4

Hence, the solution for the given quadratic equation is (-1 ± √7i) / 4.

Thanks